OpenStudy (anonymous):

What is the relationship between the number of radioactive atoms present compared with the number of daughter atoms in an element across time? After time has gone by, are their more or less radioactive atoms present? Why? Can the rate of radioactive decay be predicted precisely? Why or why not?

OpenStudy (anonymous):

Yes, radioactive decay can be predicted precisely by the below equation ... Equation for radioactive decay: N / No = e^-(0.693·t / t1/2) where, N = number of radioactive atoms present after time t No = initial number of daughter atoms N / No = fraction of radioactive atoms present after time t. t = time t1/2 = half-life The equation is precise, but if the data used isn't, then that's another story. "Garbage in, garbage out" There are less radioactive atoms present over time, because the word radiation means to 'emit or lose' particles

OpenStudy (anonymous):

lol that is from Yahoo answers